OpenStudy (anonymous):

Find the absolute max and min values of f(x) = x/x^2-x+1 on the interval [0,3]

myininaya (myininaya):

Have you found f'?

OpenStudy (anonymous):

I used quotient rule and got f' (x) = -x^2 + 1/ (x^2+1)^2

myininaya (myininaya):

\[f(x)=\frac{x}{x^2-x+1} \] \[f'(x)=\frac{(x)'(x^2-x+1)-x(x^2-x+1)'}{(x^2-x+1)^2}\] Is this what you did to find f'?

myininaya (myininaya):

(x)'=1 (x^2-x+1)'=2x-1 Plug those in and simplify the top.

myininaya (myininaya):

Then find when f' dne and find when f'=0 (think about the domain of f). If the numbers you found from f' dne and f'=0 are numbers that exist in your domain for f, then these are critical numbers for f. Plug in your critical numbers into f along with your endpoints to find the absolute highest point and absolute lowest point on your graph.

OpenStudy (anonymous):

okay I wrote the question wrong the function is f(x) = x/x^2 + 1

OpenStudy (anonymous):

so doing quotient rule f' = 1(x^2+1) - x(2x) / (x^2+1) ^2

OpenStudy (anonymous):

so simplified its -x^2 +1 / (x^2+1)^2

OpenStudy (anonymous):

now im stuck, how do i find the critical numbers? do I just look at the denominator

myininaya (myininaya):

when is -x^2+1 equal to 0? when is (-x^2+1)/(x^2+1)^2 not exist? (it exist everywhere. the bottom is never 0) So just answer the first question.

OpenStudy (anonymous):

when x = 1, or -1

myininaya (myininaya):

Ok great now your domain for your function was [0,3] and -1 is not included in that so we can throw -1 in the trash but we are keeping 1 since 1 is in [0,3]

myininaya (myininaya):

Now plug 0,1,3 into your original function

OpenStudy (anonymous):

okay

myininaya (myininaya):

find these \[f(0), f(1),f(3)\]

OpenStudy (anonymous):

f(1) is my abs max since it has the highest value

myininaya (myininaya):

ok great.

myininaya (myininaya):

now what is abs min?

OpenStudy (anonymous):

f(0)

OpenStudy (anonymous):

so quick question, for the denominator part you just look at values to see when it does not exist ? so if it were to have a restriction, then that value would also be a critical number?

myininaya (myininaya):

yep if that number is in the domain of f.

OpenStudy (anonymous):

ok thanks